Hello I wanted to ask for your help because I have an algebra problem about inequalites that am not sure what to do. Well the problem says:
To get a grade of C in your course, you must average at least 75% on 4 exams. You have taken the 1st three exams and gotten scores of 68, 78, 81. What must you score on the last exam to get a C or better?
Please answer soon. Thank you.

Answer by rfer(16322)

(68+78+81+x)/4=75
227+x=75*4
227+x=300
x=300-227
x=73%